How much 12% acid solution and pure water must be mixed to produce 228 gallons of a mixture that is 7% acid?

Answer by TimothyLamb(4379)

x = 12% acid = 88% water
y = 100% water
---
7% acid = 93% water
---
88x + 100y = 93(x + y)
x + y = 228
---
put the system of linear equations into standard form
---
88x + 100y = 93*228
x + y = 228
---
88x + 100y = 21204
x + y = 228
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x = 12% acid = 133 gal
y = 100% water = 95 gal
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php